# Label Anything: Multi-Class Few-Shot Semantic Segmentation with Visual Prompts

### v1

@@ Abstract
We present Label Anything, an innovative neural network architecture designed for few-shot semantic segmentation (FSS) that demonstrates remarkable generalizability across multiple classes with minimal examples required per class. Diverging from traditional FSS methods that predominantly rely on masks for annotating support images, Label Anything introduces varied visual prompts—points, bounding boxes, and masks—thereby enhancing the framework’s versatility and adaptability. Unique to our approach, Label Anything is engineered for end-to-end training across multi-class FSS scenarios, efficiently learning from diverse support set configurations without retraining. This approach enables a “universal” application to various FSS challenges, ranging from $1$-way $1$-shot to complex $N$-way $K$-shot configurations while remaining agnostic to the specific number of class examples. This innovative training strategy reduces computational requirements and substantially improves the model’s adaptability and generalization across diverse segmentation tasks. Our comprehensive experimental validation, particularly achieving state-of-the-art results on the COCO-$20^{i}$ benchmark, underscores Label Anything’s robust generalization and flexibility. The source code is publicly available at: https://github.com/pasqualedem/LabelAnything.
Keywords:
Few-Shot Semantic Segmentation, Visual Prompts, Multi-Class Segmentation, Class Prototypes, Deep Learning Architectures
Figure 1: Overview of the proposed framework, illustrating the workflow for a 1-shot scenario involving three classes, two support images, and all three types of prompts. Label Anything leverages a pre-trained ViT to derive image features. Prompts are fed into the Prompt Encoder, which generates a distinct prototype for each class. Subsequently, these prototypes inform the Mask Decoder, facilitating accurate pixel classification across the query image.
@@ 1 Introduction
The advancement of semantic segmentation has been closely linked to the availability of extensive labeled datasets, which often pose a significant challenge in deploying these systems for real-world applications. Typically, models trained on such datasets excel at recognizing and classifying objects they were trained on but falter when encountering new concepts or domains, thus limiting their practical utility. Segmenting objects from previously unseen classes with few or no labeled examples remains a pivotal challenge in computer vision.
@@
Few-shot semantic segmentation (FSS) has recently gained attention as the task of segmenting a query image based on a limited set of examples [26]. While the majority of research in FSS has concentrated on binary segmentation tasks (distinguishing background from foreground), there has been limited exploration of methods capable of segmenting multiple classes simultaneously, a step toward more flexible and domain-agnostic segmentation [4, 33, 14, 42, 13, 30]. Furthermore, aside from an initial exploration by Rakelly et al. [24], which considered coordinates as prompts, the field has predominantly focused on mask annotations as the sole input type. This underscores the growing need to support diverse input annotations, reducing reliance on extensive human annotation and facilitating smoother integration with other models, such as object detectors that generate bounding boxes.
@@
This work presents Label Anything (LA), a novel deep neural network architecture designed for multi-prompt, multi-class FSS. Our method embraces a prototype-based learning paradigm, leveraging diverse visual prompts—points, bounding boxes, and masks—to create a highly flexible and generalizable framework (Fig. 1 illustrates our approach). This innovation not only addresses the challenge of segmenting objects from unseen classes with limited examples but also significantly reduces the annotation burden.
@@
Building upon Segment Anything’s innovative use of diverse prompts for segmentation [9], our method significantly enriches FSS by enhancing support set adaptability and extending the framework into multi-class challenges. Unlike traditional approaches that focus on binary segmentation and adhere to strict class example limits, LA supports various prompt types and a flexible number of class examples per support set. This flexibility allows for capturing complex inter-class relationships from a single image, improving segmentation precision. Moreover, LA’s architecture is designed to learn from different support set configurations and prompt types in a single training cycle, eliminating the need for retraining across different FSS scenarios. Such versatility ensures that LA can be applied across a wide range of domains, offering a comprehensive solution that moves beyond the constraints of conventional FSS models.
Our extensive experimental analysis on the well-known COCO-$20^{i}$ benchmark [22, 33] reveals that our model achieves state-of-the-art performance, setting a new standard in the field. The contributions of LA are not merely technical but also practical as we pave the way for more accessible and efficient semantic segmentation, especially in scenarios where annotation resources are scarce.
@@
The rest of this paper is organized as follows. Section 2 reviews related work. Section 3 details our proposed approach. Section 4 evaluates our model’s performance. Section 5 concludes the paper with a summary of our findings and directions for future research.
@@ 2 Related Work
2.1 Few-Shot Semantic Segmentation
@@
Semantic segmentation, which involves assigning a class label to each pixel in an image, is a fundamental task in computer vision. The introduction of Convolutional Neural Networks (CNNs), with pioneering architectures like the Fully-Convolutional Network [17] and U-Net [25], was pivotal in pushing the field forward. These architectural innovations laid the foundation for numerous subsequent methods, each tailored to address specific challenges (e.g., [3, 23]).
@@
Building upon few-shot learning principles [6, 32, 28, 29], few-shot semantic segmentation has developed as a significant extension of semantic segmentation to achieve high accuracy with minimal reliance on large, annotated datasets. FSS involves per-pixel classification in a query image, guided by a “support set” typically comprising a small number of images and their corresponding mask annotations. This task hinges on the model’s ability to leverage support set information for accurate segmentation.
@@
Early attempts in FSS utilized CNNs to synthesize support set information. This information was integrated into the segmentation process by modifying weights and biases or merging it with query image features [26, 24]. This foundational work laid the groundwork for more sophisticated approaches broadly categorized into prototype-based and affinity-based learning. Prototype-based methods focus on deriving one or several representative prototypes for the concept to be segmented, applying these to the query image through distance metrics or decoder networks [4, 43, 33, 39, 10, 14, 15, 42]. In contrast, affinity learning approaches delve into the pixel-wise relationships between support and query features [40, 27, 35, 1, 21]. Nearly all methods in FSS leverage the extensive feature pyramids extracted by CNNs for query and support images. Only recent works such as FPTrans [41] have begun to explore the use of the Vision Transformer (ViT) [5] as the encoder.
In recent years, considerable interest has been shown in FSS. However, most research has focused on a binary paradigm, where images are segmented based on a single concept specified by the support set. The binary (single-way) and multi-class FSS settings present distinct challenges. Binary classification models struggle to adapt to multi-class scenarios without a significant drop in performance, highlighting the need for specialized solutions.
A limited number of studies have ventured into the complex domain of multi-class FSS [4, 33, 14, 42, 13, 30]. This involves enabling the support set to introduce multiple classes for segmentation, requiring the model to distinguish between $N$ predefined classes in the query image, a departure from the more straightforward foreground-background dichotomy. Multi-class FSS typically adopts an $N$-way $K$-shot framework, entailing $N$ classes within the support set and $K$ instances per class. This approach, geared towards segmenting multiple classes, invariably relies on a prototype-based model architecture to encapsulate and convey class-specific information throughout the network. For example, MFNet [42] leverages Masked Global Pooling alongside attention mechanisms to generate class prototypes that integrate with query image features for segmentation.
It is crucial to recognize that the efficacy of multi-way methodologies is contingent upon predefined $N$ and $K$ values, necessitating their specification prior to training—a constraint that also influences many single-way models in terms of the number of examples per segmented concept.
2.2 Promptable Segmentation
Integrating prompt-based learning, inspired by its success in natural language processing, into computer vision introduced a novel paradigm. In particular, a significant breakthrough came from the Segment Anything model (SAM), introduced by Kirillov et al. [9]. This innovation harnesses an interactive, prompt-guided vision framework, representing a pivotal moment reminiscent of the GPT series but in computer vision. SAM has been trained on a vast dataset comprising over 1 billion masks extracted from 11 million images, employing a “promptable” segmentation task. This innovative approach equips SAM with the unique capability for robust zero-shot generalization, enabling it to excel in a wide range of tasks such as edge detection and instance segmentation.
@@ 2.2 Promptable Segmentation
In recent times, the research community has been prolific in extending the capabilities of SAM and exploring its potential across diverse domains. These endeavors have pushed the boundaries of what SAM can accomplish and have successfully applied it to a wide range of tasks, including medical image analysis [20], image inpainting [37], and image editing [36]. Recent innovations like SegGPT [34] and SEEM [44] have extended the use of prompts to guide predictions, exploring the integration of textual and audio prompts to enhance model performance across diverse tasks.
@@
However, despite these advancements, there remains a gap in applying these models to FSS, especially in classifying objects into previously unseen classes based on a limited number of annotated examples. While models like SegGPT show promise in FSS scenarios, their generalization capabilities are often attributed to the vast datasets used for training rather than an inherent ability to leverage support set information effectively. Furthermore, apart from an early exploration by Rakelly et al. [24], research in FSS has scarcely investigated the use of diverse annotations within support sets—such as points or bounding boxes—as alternatives to traditional mask prompts, leaving a substantial area of FSS largely unexplored.
2.3 Our Contribution
Our work seeks to bridge the gaps in the literature using various visual prompts to guide the FSS task. Drawing from cutting-edge prompt-based segmentation techniques, LA innovatively employs an array of prompt types within the FSS framework. Our novel end-to-end neural network architecture is also designed for multi-class FSS, enhancing the versatility of forming support sets, crucial for effective FSS. Moving beyond the conventional focus on binary FSS, which mainly leverages affinity-based methods, our work proposes a prototype-based architecture. This design choice aligns with the efficiency and coherence demands of multi-way scenarios and addresses the limitations imposed by the significant memory requirements of affinity-based approaches in $N$-way $K$-shot setups. Taking cues from successful $1$-way $K$-shot models (e.g., [21, 31, 38, 27]), which adapt to generic $K$-shot scenarios post-training, our method expands this approach to accommodate multi-class contexts. We train our model once, enabling it to seamlessly adapt to various FSS benchmarks, including $1$-way $1$-shot and $N$-way $K$-shot configurations, without imposing any constraints on the number of examples ($K$) per class. This approach affords unparalleled flexibility, allowing the model to interpret support sets comprising any number of images and annotations, irrespective of class or prompt type. Such enhancements inherently reduce computational demands and significantly boost the model’s versatility and robustness in diverse FSS scenarios.
3 Label Anything
We aim to develop a model capable of labeling any image across any class using a minimal set of prompts for each class. This approach does not prescribe a predefined number of prompts per class; instead, it mandates that the support set comprehensively covers all required classes.
3.1 Problem Formulation
We introduce the task of multi-prompt, multi-way, few-shot semantic segmentation. This involves classifying the pixels of a query image $I_{q}$ into one of $N+1$ classes, with $N$ denoting the number of foreground categories to be segmented. The $N$ foreground concepts are learned using only a few annotated images. To achieve this goal, we employ “episodic training” [32], which consists of training episodes involving a query image and a distinct support set.
@@ 3.1 Problem Formulation
Formally, our dataset is partitioned into two subsets: a training set $\mathcal{D}_{train}$ and a test set $\mathcal{D}_{test}$. Crucially, we define two distinct sets of segmentation categories, $\mathcal{C}_{seen}$ for training and $\mathcal{C}_{unseen}$ for testing, ensuring $\mathcal{C}_{seen}\cap\mathcal{C}_{unseen}=\emptyset$. Annotations in $\mathcal{D}_{train}$ utilize classes from $\mathcal{C}_{seen}$, whereas $\mathcal{D}_{test}$ is evaluated on classes from $\mathcal{C}_{unseen}$, representing concepts not encountered during training to assess the model’s generalization capabilities in a few-shot setting.
@@
Each training episode presents the model with a query and its corresponding ground truth $(I_{q},M_{q})$, where $I_{q}$ is an RGB image of dimensions $3\times H\times W$ and $M_{q}$ is a segmentation mask associating pixels of $I_{q}$ with one of $N+1$ classes. The model receives a support set $\mathcal{S}=\{(I_{i},A_{i})\}_{i=1}^{L}$ comprising $L$ image-prompt annotation pairs to assist in classifying the query pixels. Each pair $(I_{i},A_{i})$ is called a “shot”, with $L$ usually being a small number. For instance, in multi-class FSS tasks, $L$ is often determined as $N\times K$, with $K$ indicating the number of examples, or shots, per class necessary for the model to learn about class $n$. Standard testing configurations on benchmark datasets typically involve setting $N$ and $K$ to values from the set $\{1,2,5\}$.
@@
Our approach significantly extends the flexibility of FSS by allowing for multiple types of prompts—masks, points, and bounding boxes—within the support set. Unlike previous FSS methodologies, LA does not require the support set size $L$ to match $N\times K$ strictly, enabling variable numbers of shots per class. Additionally, LA permits a single image to serve as a support example for multiple classes without repetition, linking different class prompts to the same image, thus maximizing the exploitation of inter-class relationships within annotations.
In more detail, we have $A_{i}=\{(M_{i,n},P_{i,n},B_{i,n})\}_{n=1}^{N}$ for each support set image, where $M_{i,n}$ is a binary mask for class $n$ in image $i$, $P_{i,n}$ represents point coordinates, and $B_{i,n}$ denotes bounding boxes. The prompt type for class $n$ in image $i$ is randomly selected during training, with padding applied as necessary for absent classes or prompt types. Evaluation on $\mathcal{D}_{test}$ involves support set annotations from $\mathcal{C}_{unseen}$, testing the model’s adaptability to new classes.

### v2

@@ Abstract
Few-shot semantic segmentation aims to segment objects from previously unseen classes using only a limited number of labeled examples. In this paper, we introduce Label Anything, a novel transformer-based architecture designed for multi-prompt, multi-way few-shot semantic segmentation. Our approach leverages diverse visual prompts—points, bounding boxes, and masks—to create a highly flexible and generalizable framework that significantly reduces annotation burden while maintaining high accuracy. Label Anything makes three key contributions: (i) we introduce a new task formulation that relaxes conventional few-shot segmentation constraints by supporting various types of prompts, multi-class classification, and enabling multiple prompts within a single image; (ii) we propose a novel architecture based on transformers and attention mechanisms, eliminating dependency on convolutional networks; and (iii) we design a versatile training procedure allowing our model to operate seamlessly across different $N$-way $K$-shot and prompt-type configurations with a single trained model. Our extensive experimental evaluation on the widely used COCO-$20^{i}$ benchmark demonstrates that Label Anything achieves state-of-the-art performance among existing multi-way few-shot segmentation methods, while significantly outperforming leading single-class models when evaluated in multi-class settings. Code and trained models are available at https://github.com/pasqualedem/LabelAnything.
[…]
9233
@@ 1 Introduction
Figure 1: Overview of the proposed framework, illustrating the workflow for a scenario involving three classes, two support images, and all three types of prompts. Label Anything leverages a pre-trained ViT to derive image features. Prompts are fed into the Prototype Learner, which generates a distinct prototype for each class. Subsequently, these prototypes inform the Mask Decoder, facilitating accurate pixel classification across the query image.
@@
The progress of semantic segmentation has been closely tied to the availability of extensive labeled datasets, which often pose a significant challenge in deploying these systems for real-world applications. Typically, models trained on such datasets excel at recognizing and classifying objects they were trained on but falter when encountering new concepts or domains, thus limiting their practical utility. Segmenting objects from previously unseen classes with few or no labeled examples remains a pivotal challenge in computer vision.
@@
Few-shot semantic segmentation (FSS) has gained attention as the task of segmenting a query image based on a limited set of examples [25]. While the majority of research in FSS has concentrated on binary segmentation tasks (distinguishing background from foreground), there has been limited exploration of methods capable of segmenting multiple classes simultaneously, a step toward more flexible and domain-agnostic segmentation [4, 13, 14, 29, 31, 40]. Furthermore, aside from an initial exploration by Rakelly et al. [23], which considered coordinates as prompts, the field has predominantly focused on mask annotations as the sole input type. This underscores the growing need to support diverse input annotations, reducing reliance on extensive human annotation and facilitating smoother integration with other models, such as object detectors that generate bounding boxes.
@@
This work presents Label Anything (LA), a novel deep neural network architecture designed for multi-prompt, multi-class FSS. Our method adopts a prototype-based learning paradigm, utilizing diverse visual prompts—points, bounding boxes, and masks—to establish a highly flexible and generalizable framework (Fig. 1 illustrates our approach). This innovation not only addresses the challenge of segmenting objects from unseen classes with limited examples but also significantly reduces the annotation burden. Building upon Segment Anything’s use of diverse prompts for segmentation [9], our method significantly enriches FSS by enhancing the adaptability of the support set and extending the framework to multi-class challenges.
@@
Unlike traditional approaches that focus on binary segmentation and adhere to strict class example limits, LA supports various prompt types and a flexible number of class examples per support set. This flexibility enables the capture of complex inter-class relationships from a single image, thereby improving segmentation precision. Moreover, LA’s architecture is designed to learn from different support set configurations and prompt types within a single training cycle, thereby eliminating the need for retraining across various FSS scenarios. Such versatility ensures that LA can be applied across a wide range of domains, offering a comprehensive solution that moves beyond the constraints of conventional FSS models.
Our extensive experimental analysis on the widely used COCO-$20^{i}$ benchmark [21, 31] shows that our model achieves state-of-the-art performance among existing multi-way FSS methods. Moreover, it significantly outperforms leading single-class FSS models when evaluated in multi-class settings, highlighting their limited adaptability to real-world scenarios where multiple object categories must be segmented simultaneously. These results underscore the need for dedicated multi-class FSS solutions. Beyond its technical contributions, LA offers practical advantages by enabling more accessible and efficient semantic segmentation, particularly in contexts where annotated data is scarce.
In summary, this work makes three contributions to advancing the field of semantic segmentation:
  • We introduce a new task, multi-prompt multi-way few-shot semantic segmentation, which emerges from a substantial relaxation of conventional few-shot segmentation constraints. This task enhances the applicability of segmentation models by supporting various types of prompts and multi-class classification, and enables multiple prompts within a single image, thereby overcoming the limitations of prior methods that require one image per prompt.
  • We propose a novel model architecture based on transformers and attention mechanisms. This purely transformer-based approach contrasts with previous convolution-dependent models, paving the way for more flexible and scalable segmentation solutions.
  • We design a versatile training procedure that enables our model to be trained once and operate seamlessly across various $N$-way $K$-shot and prompt-type configurations, unlike prior methods that require separate models for each parameter combination.
@@
The rest of this paper is organized as follows. Section 2 reviews related work. Section 3 introduces the task of multi-prompt multi-way FSS derived from a relaxation of the typical FSS constraints. Section 4 details our proposed approach. Section 5 evaluates our model’s performance. Section 6 concludes the paper with a summary of our findings and directions for future research.
@@ 2 Related Work
Few-Shot Semantic Segmentation
@@
Semantic segmentation, which involves assigning a class label to each pixel in an image, is a fundamental task in computer vision. The introduction of convolutional neural networks (CNNs), with pioneering architectures such as the Fully Convolutional Network [16] and U-Net [24], was pivotal in advancing the field. These architectural innovations laid the foundation for numerous subsequent methods, each tailored to address specific challenges (e.g., [3, 22]).
@@
Building upon few-shot learning principles [6, 27, 28, 30], few-shot semantic segmentation has evolved as a significant extension of semantic segmentation, achieving high accuracy with minimal reliance on large, annotated datasets. FSS involves per-pixel classification in a query image, guided by a “support set” typically comprising a small number of images and their corresponding mask annotations. This task hinges on the model’s ability to leverage support set information for accurate segmentation.
@@
Early attempts in FSS utilized CNNs to synthesize support set information. This information was integrated into the segmentation process by modifying weights and biases or merging it with query image features [23, 25]. This foundational work laid the groundwork for more sophisticated approaches broadly categorized into prototype-based and affinity-based learning. Prototype-based methods focus on deriving one or several representative prototypes for the concept to be segmented, applying these to the query image through distance metrics or decoder networks [4, 10, 14, 15, 31, 37, 40, 41]. Most existing approaches employ CNNs as feature extractors; more recently, FPTrans [39] explored the use of Vision Transformers (ViT) [5] as encoder. In contrast, affinity learning approaches delve into the pixel-wise relationships between support and query features [1, 20, 26, 34, 38]. While affinity-based approaches often achieve stronger performance, they come at a high computational cost. These methods require access to and processing of the entire support set during inference for each query image, making them challenging to scale—especially in multi-class (multi-way) settings. In fact, all existing multi-way methods are prototype-based, as they enable precomputed, reusable class representations, making them more scalable and efficient in real-world scenarios.
However, most recent developments have focused on a binary paradigm, where images are segmented based on a single concept specified by the support set. The binary (single-way) and multi-class FSS settings present distinct challenges. Binary classification models struggle to adapt to multi-class scenarios without a significant performance drop, highlighting the need for specialized solutions. As a result, only a few studies have explored the complex area of multi-class FSS [4, 13, 14, 29, 31, 40]. This involves enabling the support set to introduce multiple classes for segmentation, requiring the model to distinguish between $N$ foreground classes in the query image, leveraging a set of $K$ support images per class, solving a prototype-based $N$-way $K$-shot task. For example, MFNet [40] leverages masked global pooling alongside attention mechanisms to generate class prototypes that integrate with query image features for segmentation. To our knowledge, MFNet represents the most recent effort explicitly addressing multi-way FSS—a direction that has since been largely overlooked in the literature. Despite this, the multi-class formulation is arguably more practical and relevant to real-world applications than single-class segmentation, which considers only one object category at a time.
Nevertheless, existing methods suffer from two main limitations: (i) the performance of multi-way approaches critically depends on the predefined values of $N$ and $K$, which must be fixed prior to training—this also impacts many single-way models by constraining the number of examples per segmented concept; (ii) these methods typically treat different classes within the same image independently, repeating the same image for each class, resulting in a more complex formulation and increased computational overhead. To address these limitations, we revisit the multi-class FSS task from a more flexible and scalable perspective. By relaxing rigid assumptions and architectural constraints, we propose a more adaptable solution that can handle diverse and flexible segmentation scenarios.
Promptable Segmentation
@@
Integrating prompt-based learning, inspired by its success in natural language processing, into computer vision introduced a novel paradigm. In particular, a significant breakthrough came from the Segment Anything Model (SAM), introduced by Kirillov et al. [9]. The research community has been prolific in extending the capabilities of SAM and exploring its potential across diverse domains. These endeavors have pushed the boundaries of what SAM can accomplish and have successfully applied it to a wide range of tasks, including medical image analysis [19], image inpainting [36], and image editing [35]. Recent innovations, such as SegGPT [33] and SEEM [42], have extended the use of prompts to guide predictions, exploring the integration of textual and audio prompts to enhance model performance across diverse tasks.
@@
However, despite these advancements, a gap remains in applying these models to FSS, particularly in segmenting objects into previously unseen classes based on a limited number of annotated examples. While models like SegGPT show promise in FSS scenarios, their generalization capabilities are often attributed to the vast datasets used for training, rather than an inherent ability to leverage support set information effectively. Furthermore, apart from an early exploration by Rakelly et al. [23], research in FSS has scarcely investigated the use of diverse annotations within support sets—such as points or bounding boxes—as alternatives to traditional mask prompts, leaving a substantial area of FSS largely unexplored. For this reason, we relax the usage of the binary mask only as support information by introducing bounding boxes and points. Specifically, we design a new prototype learning stage, in which diverse prompts are encoded into a common, shared representation, thereby enhancing our model’s flexibility to accommodate various types of prompts.
3 Multi-Prompt Multi-Way FSS
We introduce the task of multi-prompt, multi-way, few-shot semantic segmentation. While the standard single-prompt, multi-way few-shot task involves classifying each pixel in a query image into one of $N$ foreground classes using $N\times K$ support image-mask pairs—with one mask per image—our approach extends this setting. Specifically, we classify the pixels of a query image $I_{q}$ into $N$ classes using $L$ examples, allowing for a variable number and type of shots per class, as well as the inclusion of multiple classes within a single support image, which is not possible in the traditional setting. Furthermore, we extend beyond the sole usage of masks as a prompt modality by introducing multiple prompt types, which can be either points or bounding boxes, to reduce human effort in shot labeling significantly.
@@ 3 Multi-Prompt Multi-Way FSS
To tackle few-shot tasks, the dataset $\mathcal{D}$ must be partitioned into two subsets: a training set $\mathcal{D}_{\text{train}}$ and a test set $\mathcal{D}_{\text{test}}$. Crucially, we define two distinct sets of segmentation categories, $\mathcal{C}_{\text{seen}}$ for training and $\mathcal{C}_{\text{unseen}}$ for testing, ensuring $\mathcal{C}_{\text{seen}}\cap\mathcal{C}_{\text{unseen}}=\emptyset$. Annotations in $\mathcal{D}_{\text{train}}$ utilize classes from $\mathcal{C}_{\text{seen}}$, whereas $\mathcal{D}_{\text{test}}$ is evaluated on classes from $\mathcal{C}_{\text{unseen}}$, representing concepts not encountered during training to assess the model’s generalization capabilities in a few-shot setting.
@@
During each training “episode”, the model receives a query image and its corresponding ground truth pair $(I_{q},M_{q})$, where $I_{q}\in\mathbb{R}^{3\times H\times W}$ represents the RGB query image and $M_{q}\in\mathbb{R}^{(N+1)\times H\times W}$ is the segmentation mask. The mask $M_{q}$ assigns each pixel in $I_{q}$ to one of $N+1$ classes, including the background. In addition, the model is provided with a support set $\mathcal{S}=\{(I_{i},A_{i})\}_{i=1}^{L}$, consisting of $L$ image-prompt annotation pairs that define the concepts to be segmented. Each pair $(I_{i},A_{i})$ is referred to as a “shot”, with $L$ typically being a small number. In multi-class FSS tasks, $L$ is commonly set as $N\times K$, where $N$ is the number of classes and $K$ is the number of examples (shots) per class.
@@
In more detail, we have $A_{i}=\{(M_{i,n},P_{i,n},B_{i,n})\}_{n=1}^{N}$, where $M_{i,n}$ is a binary mask for class $n$, $P_{i,n}$ represents point coordinates, and $B_{i,n}$ denotes bounding boxes, harnessing the multi-prompt setting. Our approach significantly extends the flexibility of FSS by allowing for multiple types of prompts within the support set. Unlike previous FSS methodologies, LA does not require the support set size $L$ to match $N\times K$ strictly, enabling variable numbers of shots per class. Additionally, LA permits a single image to serve as a support example for multiple classes without repetition, linking different class prompts to the same image, thus maximizing the exploitation of inter-class relationships within annotations, enhancing flexibility and reducing human effort in shot labeling.
4 Proposed Method
We introduce Label Anything, a model for few-shot segmentation tailored to the multi-prompt, multi-way setting. Given a support set consisting of a few images with visual prompts for each class, LA enables the semantic segmentation of user-specified target classes. The overall framework is illustrated in Fig. 1. This section details the model architecture, training procedure, and key design choices that allow LA to generalize effectively to diverse and previously unseen classes. For clarity, we omit the batch size dimension (corresponding to the episode axis) in the following explanation; it is reintroduced when describing the training procedure in Sec. 4.4.